OpenStudy (lilai3):

PLEASE HELP. WILL REWARD WITH A MEDAL OR A RIDDLE. . OR AN OREO. A certain canoe was advertised as having an "800 pound capacity," meaning that it can hold at most 800 pounds. If four people plan to use the canoe and take 60 pounds of supplies, write and solve an inequality to find the average weight per person. EXPLAIN. THANK YOU.

OpenStudy (kc_kennylau):

a riddle please

OpenStudy (kc_kennylau):

Let the average weight per person be p kg. 4(p+60)≤800

OpenStudy (kc_kennylau):

Each person holds 60 pounds of supplies, therefore p+60 There are four "persons and supplies", therefore 4(p+60) At most 800, which means smaller than or equal to 800, therefore ≤800

OpenStudy (anonymous):

i'll take a peanutbutter oreo please

OpenStudy (kc_kennylau):

i want a riddle please thank you

OpenStudy (anonymous):

I am smaller when I am low and grow as I rise, what am I?

OpenStudy (primeralph):

Wallet. Bread(dough).

OpenStudy (anonymous):

additional clue same riddle, at the bottom of the ocean I am quite small, by the time I rise to the surface I am so big that I do not exist at all.

OpenStudy (anonymous):

additional clue, all children love me

OpenStudy (lilai3):

4n+60<=800?

OpenStudy (anonymous):

I come in more colors than KC's avatar pic

OpenStudy (kc_kennylau):

4(n+60) not 4n+60, because there are 4 supplies

OpenStudy (lilai3):

oh right. thanks.

OpenStudy (anonymous):

answer: a balloon fill with compressed air at depth.

OpenStudy (lilai3):

Okay, since @kc_kennylau wanted a riddle, I will give you one. I have holes in my top and bottom, my left and right, and in the middle. But I still hold water. What am I?

OpenStudy (anonymous):

sponge

OpenStudy (kc_kennylau):

swimming pool

OpenStudy (lilai3):

sponge

OpenStudy (kc_kennylau):

lolz i suck

OpenStudy (lilai3):

wait... everyone got 60 pounds EACH, or just in total....because 60 pounds is a lot....and wouldn't the problem say "EACH" if each person had 60 pounds of luggage with them...so aren't I correct?

OpenStudy (kc_kennylau):

oh sorry

OpenStudy (kc_kennylau):

:/

OpenStudy (anonymous):

4x +60(then than/equal to) 800 x = 185

OpenStudy (anonymous):

x also = average wieght of passenger

OpenStudy (lilai3):

so am i or aren't I correct?

OpenStudy (kc_kennylau):

\[\begin{array}{rcl}4x+60&\le&800\\4x&\leq&740\\x&\le&185\end{array}\]Yep you're correct

OpenStudy (anonymous):